OpenStudy (anonymous):

what is the volume of a square Peerman with the base edges of 40 cm and a slant height of 25 cm 533.3 cm^3 8,000 cm^3 12,000 cm^3 24,000 cm^3

OpenStudy (anonymous):

\[V= \frac{ 1 }{ 3 }b ^{2}h\]This equation is just for a square pyramid and uses height (altitude). In this problem you were given the slant height (length of hypotenuse). To calculate the height use the pythagorean theorem\[a ^{2} + b ^{2} = c ^{2}\] c in this case = 25cm. a = half of 40 which is 20cm. So it can be written as 20^2 + h^2 = 25^2 400 + h^2 = 625 minus 400 from both sides h^2 = 225 take the square root from both sides h = 15 Now calculate \[\frac{ 40^{2}15 }{ 3 }\]